# Chapter 0 Strong CP problem, theta term and QCD topological properties

### v1

@@ 3 General aspects of $\theta$-dependence
As we argued above, within the general framework of the Standard Model one cannot exclude a priori a $\theta$ term in the QCD Lagrangian. At the same time, the dependence of various properties of strong interactions on $\theta$ plays a significant role by itself, which is relevant to many phenomenological and theoretical issues, like the interpretation of some particular features of the hadronic spectrum, or the characterization of the various phases which are found when considering Yang–Mills theories or QCD at finite temperature. Moreover, one of the proposed solutions to the strong CP problem is the so-called Peccei–Quinn mechanism Peccei:1977hh ; Peccei:1977ur , which invokes the presence of a new particle, the axion, which has also become a popular candidate for dark matter: since the axion couples to QCD like a $\theta$ term, the study of $\theta$-dependence in QCD turns out to be fundamental in order to predict some properties of this dark matter candidate.
@@
It is clear from Eq. (38) that, in the Euclidean formulation, the $\theta$ term gives a purely imaginary contribution to the action, making it complex. This is usually known as then sign problem, since it ruins the positivity of the Euclidean path-integral measure, thus undermining the Monte Carlo approach at its very basis. A similar problem appears in other cases, for instance, the fermion determinants themselves becomes complex after adding chemical potentials coupled to quark numbers. In this case, its origin can be traced back to the fact that the $\theta$ term involves an odd number of time derivatives (contrary to the standard pure gauge term), hence the appearance of the extra factor $\mathrm{i}$”.
@@
where $\mathcal{V}$ is the 4-dimensional volume. The extra factor $\mathrm{i}$” of the Euclidean $\theta$ term implies that the ratio of partion functions can be viewed as the expectation value of a pure phase factor, $Z(T,\theta)/Z(T,0)=\braket{\mathrm{e}^{\mathrm{i}\theta Q}}_{\theta=0}$, which is then bounded by 1 from above, meaning that $F(T,\theta)\geq 0$ with a minimum in $\theta=0$ [^1] The same property holds for the quantum particle moving on a circumference discussed above: in that case, since a non-zero $\theta$ can be interpreted as the presence of a magnetic flux, the property implies that the system is diamagnetic. That rules out the possibility of spontaneous P breaking at $\theta=0$ Vafa:1984xg , hence $F(T,\theta)$ is an even function of $\theta$ with period $2\pi$ (because $Q\in\mathbb{Z}$) and analytic around $\theta=0$. Therefore, $F(T,\theta)$ can be written as a Taylor series around $\theta=0$, containing only even powers. A common parametrization is the following:
@@
$\chi=\frac{1}{\mathcal{V}}\braket{Q^{2}}\ ;\ \ \ \ \ b_{2n}=\frac{2(-1)^{n}}{(2n+2)!}\frac{\braket{Q^{2n}}_{c}}{\braket{Q^{2}}}$ (42)
@@
As we have discussed above, the problem why $\theta$ is so small does not find a solution within the Standard Model, given that the possibility of a massless up quark is ruled out. An attractive solution is provided by the so-called Peccei–Quinn mechanism Peccei:1977hh ; Peccei:1977ur ; Wilczek:1977pj ; Weinberg:1977ma , which invokes an extension of the Standard Model with an extra axial U(1) global symmetry, the U(1)PQ Peccei–Quinn symmetry. Such symmetry is supposed to be spontaneously broken, implying the possible emergence of a massless Nambu-Goldstone boson, the axion, but also anomalously broken at the same time, leading to a coupling of the axion field $a(x)$ to the topological charge density operator $q(x)$.
One can built various theories implementing such features, see e. g., Kim:1979if ; Shifman:1979if ; Dine:1981rt ; Zhitnitsky:1980tq for early proposals and DiLuzio:2020wdo for a recent review, however, at the level of a low energy effective theory, they all correspond to the addition to the QCD Lagrangian of a massless pseudoscalar with derivative couplings plus a modification of the $\theta$ term as follows:
@@ 2 Low temperature: large-$N$ and chiral scaling
It is useful to stress that this large-$N$ prediction is not necessarily at odds with the DIGA model: indeed, they are in perfect agreement if $\chi=0$, meaning no $\theta$-dependence at all. However, if $\chi\sim\mathcal{O}(N^{0})$ as required by the Witten–Veneziano mechanism, the large-$N$ argument leads to a non-trivial prediction, which differs from DIGA substantially. In this case, indeed, by matching the largeN-N limit for the scaling function, hence finite coefficients for the expansion in terms of $\bar{\theta}$, one can infer the following scaling behavior for the $b_{2n}$ coefficients:
@@ 3 The Lattice QCD approach: basics and results
The first attempts DiVecchia:1981aev ; DiVecchia:1981hh were based on a simple gluonic discretization of the topological charge density operator, representing $F_{\mu\nu}$ in terms of an open plaquette, then subtracting an additive renormalization for $\chi$ based on perturbative computations. Results provided a non-zero value for $\chi$ in the SU(2) pure gauge theory, however about 2 orders of magnitude smaller than predicted by the WV formula. Later studies clarified that the problem was in a missing multiplicative renormalization constant for the discretized topological charge density Campostrini:1988cy ; Campostrini:1989dh : after taking that into account, later determinations led to results confirming the WV mechanism also quantitatively, both for SU(2) and SU(3) pure gauge theories Alles:1996nm ; Alles:1997qe . The value of $\chi$ for pure gauge SU(3) theory has been verified by many other studies, adopting gluonic definitions in combination with smoothing techniques Durr:2006ky ; Ce:2015qha ; Bonati:2015sqt ; Athenodorou:2020ani ; Athenodorou:2021qvs ; Bonanno:2023ple ; Durr:2025qtq , or fermionic definitions DelDebbio:2004ns ; Luscher:2010ik ; Cichy:2015jra ; Bonanno:2019xhg , achieving an increasing control over the continuum extrapolation and other possible systematic effects. Recently, also the so-called slope of the topological susceptibility, related to the leading order momentum dependence of the two point function of $q(x)$, has been determined Bonanno:2023ple .
@@
For this reason, early determinations of $b_{2}$ for the SU(3) pure gauge theory were affected by statistical errors of the order of 20% or larger DelDebbio:2002xa ; DElia:2003zne ; Giusti:2007tu . A substantial improvement has been achieved by adding to the action an explicit source term coupled to the topological charge, in particular a purely imaginary $\theta$ term, so as to avoid the sign problem. In this way, the information about the $b_{2n}$ coefficients is transferred to the dependence on the source term of lower order cumulants, with a sizable reduction of statistical errors Panagopoulos:2011rb ; Bonati:2015sqt , bringing the uncertainty for SU(3) well below the 10% level. Further studies have looked at $b_{2}$ also for larger values of $N$ Bonati:2016tvi ; Bonanno:2020hht , obtaining results well compatible with the predicted $1/N^{2}$ scaling: an example, taken from Ref. Bonanno:2020hht , is reported in the right-hand side of Fig. 1. For the same reason, the determination of the other $b_{2n}$ coefficients turns out to be more and more challenging as $n$ increases: as a matter of fact, only $b_{4}$ has been determined, and only for the SU(2) pure gauge theory Bonanno:2018xtd .
@@ 4 High temperature: $\theta$-dependence across the phase transition
Early indications for a suppression of $\chi$ across the deconfinement transition of the SU(2) pure gauge theory Teper:1985gi ; Teper:1985ek were later confirmed both for SU(3) Alles:1996nm and SU(2) Alles:1997qe . In particular, the topological susceptibility appears to be practically independent of $T$ as long as $T<T_{\rm c}$, where $T_{\rm c}$ is the deconfinement transition, then rapidly dropping to small but non-zero values for $T>T_{\rm c}$. The extension of such results to larger values of $N$ DelDebbio:2004vxo then showed that $\chi$ seems indeed to vanish, as $N\to\infty$, for any $T>T_{\rm c}$, as expected from the argument above. Recently, this has been further corroborated by a calculation of the susceptibility at the critical temperature in the deconfinement phase $\chi_{\rm d}^{+}=\chi(T=T_{\rm c}^{+})$ Borsanyi:2022fub ; Bonanno:2023hhp , showing that this quantity vanishes as $\sim\exp(-N)$ approaching the large-$N$ limit Bonanno:2023hhp . This is a strong indication that the DIGA regime indeed sets in, also for finite $N$, as soon as the system crosses deconfinement: further studies of pure gauge topology at finite $T$ have further refined this picture Berkowitz:2015aua ; Borsanyi:2015cka ; Frison:2016vuc ; Kitano:2015fla ; Borsanyi:2021gqg looking also for the asymptotic behavior of $\chi$ as a function of $T$ predicted at the 1-loop level, see Eq. (47), which, as we have stressed above, could set in at temperatures much larger than those at which the diluteness hypothesis starts to be valid.
@@
When considering lattice investigations of $b_{2}$ above $T_{\rm c}$, one should actually be careful about the following possible caveat. Since $\chi$ drops rapidly above $T_{\rm c}$, and since lattice simulations are always constrained to finite volumes, it may easily happen that at high enough temperatures (or too small volumes) $\langle Q^{2}\rangle=\mathcal{V}\,\chi\ll 1$. In this situation, the gauge ensemble is essentially composed by configurations with $Q=0$, with only a small fraction $\epsilon$ with $|Q|=1$ and practically none with $|Q|>1$, so that $Z(T,\theta)\propto(1+\epsilon\cos\theta)$ and $F(T,\theta)\propto(1-\cos\theta)$. However, this is clearly a fake DIGA behaviour, driven by finite volume effects, which is quite different from a fair DIGA behaviour in which $\langle Q^{2}\rangle\gtrsim 1$ and nevertheless one has $F(T,\theta)\propto(1-\cos\theta)$ because of the Poissonian distribution of integer charged topological excitations. The lattice determinations of $b_{2}$ across $T_{\rm c}$ are typically not affected by this possible caveat.
@@ 4 Conclusions
In this chapter we introduced the $\theta$-dependence of four dimensional gauge theories and some of its main features. In particular, we discussed several approximation schemes that can be used to get quantitative or, in some cases, semi-quantitative information on the $\theta$-dependence of QCD and QCD-like theories in various regimes. We then presented a short review of recent results obtained by carrying out numerical simulations of the lattice discretized theory.
While $\theta$-dependence provides fundamental insights in several aspects of the low-energy QCD phenomenology, and especially for what concerns the pseudoscalar meson spectrum, it also opens new problems. The most famous of these problems is the strong CP problem, which is related to the experimental absence of CP violations in the strong sector of the standard model, despite the fact that the $\theta$ term explicitly breaks CP. This fact forces the physical value of the $\theta$ angle to be surprisingly small, $|\theta|\lesssim 10^{-10}$.
The axion solution of the strong CP problem is definitely the most investigated one, mainly because of the phenomenological appeal of the axion as a dark matter candidate. Since several axion properties (in particular its mass and self-coupling) can be directly related to $\theta$-dependence, in the last years there has been a renewed interest in the $\theta$-dependence of QCD at high temperature. It is indeed possible to use the high-temperature behavior of the topological susceptibility $\chi(T)$ (more precisely: $\chi(T)$ in the range of temperatures between $\approx$150MeV and OO(10)GeV) to put an upper bound on the axion decay constant. Despite the enormous progress made in recent years, several important questions regarding $\theta$-dependence remain open or at least need to be explored further.
@@
When discussing lattice result in Sec. 3 we focused on $\theta$ dependence for small values of $\theta$ and at equilibrium, discussing aspects that are, in practice, related to the second and to the fourth derivative of the free energy density $f(T,\theta)$ at $\theta=0$. In this regime, although some delicate points still need to be further investigated, we have an overall rather clear physical picture. Concerning what happens for larger values of $\theta$, and in particular at $\theta=\pi$, there are instead only exploratory studies Kitano:2017jng ; Kitano:2020mfk ; Hirasawa:2024fjt with more reliable results existing only for systems in reduced dimensionality Alles:2007br ; Alles:2014tta ; Sulejmanpasic:2020lyq ; Sugeno:2025exv ; Matsumoto:2025fjb . Direct simulations at real $\theta$ values are prevented by the sign problem, and for large $\theta$ it is difficult to obtain reliable results from imaginary $\theta$ simulations or Taylor expansion at $\theta=0$, due to the difficulties of performing the analytic continuation or to evaluate a large enough number of coefficients of the expansion. Real time results, relevant to describe out of equilibrium processes, also present huge difficulties, mainly (but not only) related to the resolution of the inverse problem required to extract spectral densities. The only results available so far in this context are those regarding the sphaleron rate Kotov:2018vyl ; Altenkort:2020axj ; BarrosoMancha:2022mbj ; Bonanno:2023ljc ; Bonanno:2023thi .
@@
[Acknowledgments] The work of C. Bonanno is supported by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and, partially, by the grant PID2021-127526NB-I00, both funded by MCIN/AEI/10.13039/501100011033. This work is also partially supported by the project “Non-perturbative aspects of fundamental interactions, in the Standard Model and beyond” funded by MUR, Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN), Bando 2022, grant 2022TJFCYB (CUP I53D23001440006).

### v2

@@ 3 General aspects of $\theta$-dependence
As we argued above, within the general framework of the Standard Model one cannot exclude a priori a $\theta$-term in the QCD Lagrangian. At the same time, the dependence of various properties of strong interactions on $\theta$ plays a significant role by itself, which is relevant to many phenomenological and theoretical issues, like the interpretation of some particular features of the hadronic spectrum, or the characterization of the various phases which are found when considering Yang–Mills theories or QCD at finite temperature. Moreover, one of the proposed solutions to the strong CP problem is the so-called Peccei–Quinn mechanism Peccei:1977hh ; Peccei:1977ur , which invokes the presence of a new particle, the axion, which has also become a popular candidate for dark matter: since the axion couples to QCD like a $\theta$-term, the study of $\theta$-dependence in QCD turns out to be fundamental in order to predict some properties of this dark matter candidate.
@@
It is clear from Eq. (38) that, in the Euclidean formulation, the $\theta$-term gives a purely imaginary contribution to the action, making it complex. This is usually known as then sign problem, since it ruins the positivity of the Euclidean path-integral measure, thus undermining the Monte Carlo approach at its very basis. A similar problem appears in other cases, for instance, the fermion determinants themselves becomes complex after adding chemical potentials coupled to quark numbers. In this case, its origin can be traced back to the fact that the $\theta$-term involves an odd number of time derivatives (contrary to the standard pure gauge term), hence the appearance of the extra factor $\mathrm{i}$”.
@@
where $\mathcal{V}$ is the four-dimensional volume. The extra factor $\mathrm{i}$” of the Euclidean $\theta$-term implies that the ratio of partion functions can be viewed as the expectation value of a pure phase factor, $Z(T,\theta)/Z(T,0)=\braket{\mathrm{e}^{\mathrm{i}\theta Q}}_{\theta=0}$, which is then bounded by 1 from above, meaning that $F(T,\theta)\geq 0$ with a minimum in $\theta=0$ [^1] The same property holds for the quantum particle moving on a circumference discussed above: in that case, since a non-zero $\theta$ can be interpreted as the presence of a magnetic flux, the property implies that the system is diamagnetic. That rules out the possibility of spontaneous P breaking at $\theta=0$ Vafa:1984xg , hence $F(T,\theta)$ is an even function of $\theta$ with period $2\pi$ (because $Q\in\mathbb{Z}$) and analytic around $\theta=0$. Therefore, $F(T,\theta)$ can be written as a Taylor series around $\theta=0$, containing only even powers. A common parametrization is the following:
@@
$\chi=\frac{1}{\mathcal{V}}\braket{Q^{2}}\ ;\ \ \ \ \ b_{2n}=\frac{2(-1)^{n}}{(2n+2)!}\frac{\braket{Q^{2n}}_{c}}{\braket{Q^{2}}}\ ,$ (42)
@@
As we have discussed above, the problem why $\theta$ is so small does not find a solution within the Standard Model, given that the possibility of a massless up quark is ruled out. An attractive solution is provided by the so-called Peccei–Quinn mechanism Peccei:1977hh ; Peccei:1977ur ; Wilczek:1977pj ; Weinberg:1977ma , which invokes an extension of the Standard Model with an extra axial U(1) global symmetry, the U(1)PQ Peccei–Quinn symmetry. Such symmetry is supposed to be spontaneously broken, implying the possible emergence of a massless Nambu–Goldstone boson, the axion, but also anomalously broken at the same time, leading to a coupling of the axion field $a(x)$ to the topological charge density operator $q(x)$.
One can built various theories implementing such features, see e. g., Kim:1979if ; Shifman:1979if ; Dine:1981rt ; Zhitnitsky:1980tq for early proposals and DiLuzio:2020wdo for a recent review, however, at the level of a low energy effective theory, they all correspond to the addition to the QCD Lagrangian of a massless pseudoscalar with derivative couplings plus a modification of the $\theta$-term as follows:
@@ 2 Low temperature: large-$N$ and chiral scaling
It is useful to stress that this large-$N$ prediction is not necessarily at odds with the DIGA model: indeed, they are in perfect agreement if $\chi=0$, meaning no $\theta$-dependence at all. However, if $\chi\sim\mathcal{O}(N^{0})$ as required by the Witten–Veneziano mechanism, the large-$N$ argument leads to a non-trivial prediction, which differs from DIGA substantially. In this case, indeed, by matching the largeN-N limit for the scaling function, hence finite coefficients for the expansion in terms of $\bar{\theta}$, one can infer the following scaling behavior for the $b_{2n}$ coefficients (for a large-$N$ analysis of QCD $\theta$-dependence using effective chiral Lagrangians see Vonk:2019kwv ; GomezNicola:2019myi ):
@@ 3 The Lattice QCD approach: basics and results
The first attempts DiVecchia:1981aev ; DiVecchia:1981hh were based on a simple gluonic discretization of the topological charge density operator, representing $F_{\mu\nu}$ in terms of an open plaquette, then subtracting an additive renormalization for $\chi$ based on perturbative computations. Results provided a non-zero value for $\chi$ in the SU(2) pure gauge theory, however about 2 orders of magnitude smaller than predicted by the WV formula. Later studies clarified that the problem was in a missing multiplicative renormalization constant for the discretized topological charge density Campostrini:1988cy ; Campostrini:1989dh : after taking that into account, later determinations led to results confirming the WV mechanism also quantitatively, both for SU(2) and SU(3) pure gauge theories Alles:1996nm ; Alles:1997qe . The value of $\chi$ for pure gauge SU(3) theory has been verified by many other studies, adopting gluonic definitions in combination with smoothing techniques Durr:2006ky ; Ce:2015qha ; Bonati:2015sqt ; Athenodorou:2020ani ; Athenodorou:2021qvs ; Bonanno:2023ple ; Durr:2025qtq , or fermionic definitions DelDebbio:2004ns ; Luscher:2010ik ; Cichy:2015jra ; Bonanno:2019xhg , achieving an increasing control over the continuum extrapolation and other possible systematic effects. Recently, also the so-called slope of the topological susceptibility, related to the leading-order momentum-dependence of the two-point function of $q(x)$, has been determined Bonanno:2023ple .
@@
For this reason, early determinations of $b_{2}$ for the SU(3) pure gauge theory were affected by statistical errors of the order of 20% or larger DelDebbio:2002xa ; DElia:2003zne ; Giusti:2007tu . A substantial improvement has been achieved by adding to the action an explicit source term coupled to the topological charge, in particular a purely imaginary $\theta$-term, so as to avoid the sign problem. In this way, the information about the $b_{2n}$ coefficients is transferred to the dependence on the source term of lower order cumulants, with a sizable reduction of statistical errors Panagopoulos:2011rb ; Bonati:2015sqt , bringing the uncertainty for SU(3) well below the 10% level. Further studies have looked at $b_{2}$ also for larger values of $N$ Bonati:2016tvi ; Bonanno:2020hht , obtaining results well compatible with the predicted $1/N^{2}$ scaling: an example, taken from Ref. Bonanno:2020hht , is reported in the right-hand side of Fig. 1. For the same reason, the determination of the other $b_{2n}$ coefficients turns out to be more and more challenging as $n$ increases: as a matter of fact, only $b_{4}$ has been determined, and only for the SU(2) pure gauge theory Bonanno:2018xtd .
@@ 4 High temperature: $\theta$-dependence across the phase transition
Early indications for a suppression of $\chi$ across the deconfinement transition of the SU(2) pure gauge theory Teper:1985gi ; Teper:1985ek were later confirmed both for SU(3) Alles:1996nm and SU(2) Alles:1997qe . In particular, the topological susceptibility appears to be practically independent of $T$ as long as $T<T_{\rm c}$, where $T_{\rm c}$ is the deconfinement transition, then rapidly dropping to small but non-zero values for $T>T_{\rm c}$. The extension of such results to larger values of $N$ DelDebbio:2004vxo then showed that $\chi$ seems indeed to vanish, as $N\to\infty$, for any $T>T_{\rm c}$, as expected from the argument above. Recently, this has been further corroborated by a calculation of the susceptibility at the critical temperature in the deconfined phase $\chi_{\rm d}^{+}=\chi(T=T_{\rm c}^{+})$ Borsanyi:2022fub ; Bonanno:2023hhp , showing that this quantity vanishes as $\sim\exp(-N)$ approaching the large-$N$ limit Bonanno:2023hhp . This is a strong indication that the DIGA regime indeed sets in, also for finite $N$, as soon as the system crosses deconfinement: further studies of pure gauge topology at finite $T$ have further refined this picture Berkowitz:2015aua ; Borsanyi:2015cka ; Frison:2016vuc ; Kitano:2015fla ; Borsanyi:2021gqg looking also for the asymptotic behavior of $\chi$ as a function of $T$ predicted at the 1-loop level, see Eq. (47), which, as we have stressed above, could set in at temperatures much larger than those at which the diluteness hypothesis starts to be valid.
@@
When considering lattice investigations of $b_{2}$ above $T_{\rm c}$, one should actually be careful about the following possible caveat. Since $\chi$ drops rapidly above $T_{\rm c}$, and since lattice simulations are always constrained to finite volumes, it may easily happen that at high enough temperatures (or too small volumes) $\langle Q^{2}\rangle=\mathcal{V}\,\chi\ll 1$. In this situation, the gauge ensemble is essentially composed by configurations with $Q=0$, with only a small fraction $\epsilon$ with $|Q|=1$ and practically none with $|Q|>1$, so that $Z(T,\theta)\propto(1+\epsilon\cos\theta)$ and $F(T,\theta)\propto(1-\cos\theta)$. However, this is clearly a fake DIGA behaviour, driven by finite volume effects, which is quite different from a fair DIGA behaviour in which $\langle Q^{2}\rangle\gtrsim 1$ and nevertheless one has $F(T,\theta)\propto(1-\cos\theta)$ because of the Poissonian distribution of integer-charged topological excitations. The lattice determinations of $b_{2}$ across $T_{\rm c}$ are typically not affected by this possible caveat.
@@ 4 Conclusions
In this chapter we introduced the $\theta$-dependence of four-dimensional gauge theories and some of its main features. In particular, we discussed several approximation schemes that can be used to get quantitative or, in some cases, semi-quantitative information on the $\theta$-dependence of QCD and QCD-like theories in various regimes. We then presented a short review of recent results obtained by carrying out numerical simulations of the lattice discretized theory.
While $\theta$-dependence provides fundamental insights in several aspects of the low-energy QCD phenomenology, and especially for what concerns the pseudoscalar meson spectrum, it also opens new problems. The most famous of these problems is the strong CP problem, which is related to the experimental absence of CP violations in the strong sector of the Standard Model, despite the fact that the $\theta$-term explicitly breaks CP. This fact forces the physical value of the $\theta$ angle to be surprisingly small, $|\theta|\lesssim 10^{-10}$.
The axion solution of the strong CP problem is definitely the most investigated one, mainly because of the phenomenological appeal of the axion as a dark matter candidate. Since several axion properties (in particular its mass and self-coupling) can be directly related to $\theta$-dependence, in the last years there has been a renewed interest in the $\theta$-dependence of QCD at high temperature. It is indeed possible to use the high-temperature behavior of the topological susceptibility $\chi(T)$ (more precisely: $\chi(T)$ in the range of temperatures between $\approx 150$ MeV and $O(10)$ GeV) to put an upper bound on the axion decay constant. Despite the enormous progress made in recent years, several important questions regarding $\theta$-dependence remain open or at least need to be explored further.
@@
When discussing lattice result in Sec. 3 we focused on $\theta$-dependence for small values of $\theta$ and at equilibrium, discussing aspects that are, in practice, related to the second and to the fourth derivative of the free energy density $f(T,\theta)$ at $\theta=0$. In this regime, although some delicate points still need to be further investigated, we have an overall rather clear physical picture. Concerning what happens for larger values of $\theta$, and in particular at $\theta=\pi$, there are instead only exploratory studies Kitano:2017jng ; Kitano:2020mfk ; Hirasawa:2024fjt with more reliable results existing only for systems in reduced dimensionality Alles:2007br ; Alles:2014tta ; Sulejmanpasic:2020lyq ; Sugeno:2025exv ; Matsumoto:2025fjb . Direct simulations at real $\theta$ values are prevented by the sign problem, and for large $\theta$ it is difficult to obtain reliable results from imaginary-$\theta$ simulations or Taylor expansion at $\theta=0$, due to the difficulties of performing the analytic continuation or to evaluate a large enough number of coefficients of the expansion. Real time results, relevant to describe out of equilibrium processes, also present huge difficulties, mainly (but not only) related to the resolution of the inverse problem required to extract spectral densities. The only results available so far in this context are those regarding the sphaleron rate Kotov:2018vyl ; Altenkort:2020axj ; BarrosoMancha:2022mbj ; Bonanno:2023ljc ; Bonanno:2023thi .
@@
[Acknowledgments] We would like to thank B. Garbrecht and F.-K. Guo for comments. The work of C. Bonanno is supported by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and, partially, by the grant PID2021-127526NB-I00, both funded by MCIN/AEI/10.13039/501100011033. This work is also partially supported by the project “Non-perturbative aspects of fundamental interactions, in the Standard Model and beyond” funded by MUR, Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN), Bando 2022, grant 2022TJFCYB (CUP I53D23001440006).